Fermi National Laboratory
 

Accelerator Update for October 20, 2006 - October 23, 2006

 

Friday October 20

The day shift began with the Tevatron (TeV) magnet replacement work continuing, with Antiproton Source (Pbar) searching for a magnet ground fault, with the Recycler, Linac, Booster, and the Main Injector (MI) operating normally, with MiniBooNE and NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) in access.

A Pbar expert found a problem at 8:32 AM with the bottom half of a quadrupole magnet (D6Q16). He had the magnet fixed by 9:55 AM.

At 11:32 AM, a water expert reported that there was a LCW leak somewhere in the TeV F sector. Operators prepared the area for an access. This access turned off Booster, so Booster experts will access their tunnel to investigate a kicker problem.

The Cryo System Run Coordinator reported at 11:35 AM that they were ready to begin the cooldown of the D1 cryo system.

At 1:08 PM, Booster experts reported that they were replacing the MKS07 kicker magnet.

At 1:54 PM, Booster experts reported that they had replaced MKS07, but MKS08 began sparking. Their investigations continued.

Operations reported 2:02 PM that everyone required has signed off on the Pbar cryo system. Experts began the cooldown.

The F sector access ended. Operations began turning Pbar and MI on at 3:47 PM.

At 3:53 PM, Operations received a fire alarm for a Pbar service building. A capacitor bank for D:IB shorted out. Experts began their repairs.

Operations established beam to Pbar and Pbar began stacking around 7:35 PM.

Saturday October 21

The midnight shift began with the evening shift operators finishing their interlock of the TeV, and with NuMI and MiniBooNE taking beam.

Operations reported at 12:05 AM that the TeV was secured.

At 1:04 AM, Operations reported a problem with the H- Source current. Operators accessed the Preaccelerator to investigate.

At 1:54 AM, Operations contacted Preaccelerator experts.

A Linac expert changed a vacuum gauge controller that was inhibiting the H- Source. An expert came in around 5 AM to tune up the source.

Operations prepared the TeV for access at 6:26 AM so that cryo technicians could work.

At 7:33 AM, Operations reported that an Accumulator device (A:IB) had tripped off. Operators called in power supply experts.

At 9:12 AM, the power supply expert investigating A:IB said there was a ground fault in the power supply cable going to the transformer. He had the cable replaced by 11:42 AM, but then he found a problem with D:QD. Operations contacted the Run Coordinator.

At 12:22 PM, power supply experts accessed the Pbar tunnel. At 12:55 PM, they reported that they had hipotted the magnet and apparently burned off the ground fault.

Pbar began stacking at 1:37 PM.

At 2:20 PM, D:QD tripped off. It reset okay, but Operations contacted an expert. At 2:46 PM Operations turned off D:QD for 15 minutes per expert instruction and then resumed stacking.

At 3:40 PM, the MI-60 MIPS tripped on a transformer oil temperature indication. Operations called in an expert who managed to reset the VCB. Pbar resumed stacking at 5:01 PM.

Operations began preparing MI at 5:17 PM for a Pbar/Recycler stash of antiprotons.

The Recycler completed their stash at 5:41 PM. Pbar resumed stacking.

Tevatron safety system testing ended at 6:10 PM.

At 6:28 PM, the MI RF vacuum suffered a vacuum burst. Operations contacted experts.

Operations reported around 7 PM that the Tevatron separators were on and in conditioning mode.

At 7:17 PM, Operations prepared MI for an access to check the vacuum. The gauges showed the vacuum at atmosphere.

At 7:35 PM, Operations reported that the MI RF vacuum valves had not closed as they should have.

Vacuum experts reported at 10:27 PM that three of the five sections were back under vacuum. At 11:51 PM, they reported that they had found and fixed a loose flange on a beam valve.

Sunday October 22

The midnight shift began with the TeV separators in conditioning mode and in cryo cooldown, and with NuMI and MiniBooNE off due to a MI vacuum problem and the access.

MiniBooNE resumed taking beam at 1:09 AM.

Operations established beam to NuMI and Pbar at 1:51 AM.

The Recycler completed a stash at 7:43 AM.

Operations reported at 8:41 AM that the TeV sector D3 cooldown had completed. Operations continued with the TeV turn on.

Operations turned of beam to NuMI at 9:44 AM per the experiment’s request. They came back on at 11:02 AM.

Operations reported there was QBS current on the TeV bus at 11:41 AM.

Operations prepared D0 and CDF for access at 2:27 PM.

At 6:08 PM, Operations reported a problem with the TeV RF. A Controls expert investigated and at 8:27 PM reported that the control cards were in the wrong mode.

Operations reported at 8 PM the Pbar stack was at 15.6E10 and the Recycler stash was at 52.2E10.

Operations turned on the TeV RF and at 9:14 PM put the TeV through a dry squeeze.

At 9:56 PM, Operations reported problems with the Tevatron Electron Lens (TEL1). Operations called in an expert.

The Recycler completed a stash at 11:08 PM.

The TEL expert reported at 11:36 PM that he was worried about the vacuum and that Operations should monitor it.

Monday October 23

The midnight shift began with the TeV turn on in progress, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

TeV experts began studies around 12:43 AM.

Operations reported at 1:26 AM that they had beam circulating in the TeV on the first try.

Operations turned on the Pelletron at 2:53 AM to prove that it was working.

The Recycler completed a stash at 4:08 AM.

At 6:15 AM, Operations reported a problem with TeV collimator readbacks. An expert rebooted the collimator front-end and replaced a control card. He had the readbacks repaired by 8:12 AM.

The Plans for Monday

The plan is for TeV studies to continue through the morning, to stack and stash, and then to establish a store by later this afternoon.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/23/2006    email Fermilab